40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 10:46:51 2010, Koji, Update, ABSL, PLL reconstructed 
    Reply  Thu Mar 18 12:27:56 2010, Alberto, Update, ABSL, PLL reconstructed 
       Reply  Thu Mar 18 13:40:35 2010, Koji, Update, ABSL, PLL reconstructed 
          Reply  Thu Mar 18 21:42:26 2010, Koji, Update, ABSL, PLL reconstructed 
             Reply  Mon Mar 22 22:11:26 2010, Mott, Update, ABSL, PLL reconstructed IMG_0150.JPG
                Reply  Mon Mar 22 23:37:32 2010, Mott, Update, ABSL, PLL reconstructed 2010-03-22_23.14.00.jpg2010-03-22_23.24.50.jpg
                   Reply  Tue Mar 23 18:44:46 2010, Mott, Update, ABSL, PLL reconstructed NPRO_PLL_freqresp.png
                      Reply  Wed Mar 24 03:58:18 2010, kiwamu, matt, koji, Update, Green Locking, leave PLL locked DSC_1402.JPG
                         Reply  Wed Mar 24 04:22:51 2010, kiwamu, matt, Update, Green Locking, two NPRO PLL  
                            Reply  Wed Mar 24 14:52:02 2010, Alberto, Update, Green Locking, two NPRO PLL  
                         Reply  Wed Mar 24 15:59:59 2010, kiwamu, matt, Update, Green Locking, leave PLL locked 
                            Reply  Thu Mar 25 17:29:48 2010, kiwamu, mott, Update, Green Locking, PLL two NPROs 
Message ID: 2714     Entry time: Thu Mar 25 17:29:48 2010     In reply to: 2712
Author: kiwamu, mott 
Type: Update 
Category: Green Locking 
Subject: PLL two NPROs 

In this afternoon, Mott and I tried to find a beat note between two NPROs which are going to be set onto each end table for green locking.

At first time we could not find any beats. However Koji found that the current of innolight NPRO was set to half of the nominal.

Then we increased the current to the nominal of 2A, finally we succeeded in finding a beat note.

Now we are trying to lock the PLL.

P.S. we also succeeded in acquiring the lock

 

nominal setup
  innolight  lightwave
T [deg] 39.75 37.27
current [A] 2 2
laser power [mW] 950 700


ELOG V3.1.3-